### v1

# Quantitative periodic homogenization of parabolic equations with large drift and potential

Abstract
This work aims to study the rates in the context of periodic homogenization of parabolic problems with large lower order terms (both drift and potential). We demonstrate that the solution is a product of three terms: (i) a function of time, (ii) the ground-state of an exponential cell eigenvalue problem and (iii) the solution to a parabolic equation with zero effective drift. For the latter, we derive $\mathrm{L}^{2}$ rates in the homogenization limit.
1 Introduction
In this paper, for the unknown $u_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, we study the following parabolic initial boundary value problem:
$\left\{\begin{aligned} \frac{\partial u_{\varepsilon}}{\partial t}-\operatorname{div}\left(A\left(\frac{x}{\varepsilon}\right)\nabla u_{\varepsilon}\right)+\frac{1}{\varepsilon}b\left(\frac{x}{\varepsilon}\right)\cdot\nabla u_{\varepsilon}+\frac{1}{\varepsilon^{2}}c\left(\frac{x}{\varepsilon}\right)u_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ u_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ u_{\varepsilon}(0,x)&=u_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (1.1)
Here the parameter $\varepsilon\in(0,1)$. The bounded domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be open and with Lipschitz boundary $\partial\Omega$. The coefficients $A,b,c$ are assumed to be bounded in $\mathrm{L}^{\infty}$ and to be $Y$-periodic with the standard notation $Y:=[0,1)^{d}$ for the unit cell. Furthermore, $A$ is assumed to be uniformly elliptic, i.e.
$A(y)\xi\cdot\xi\geq\mu\left|\xi\right|^{2}\quad\forall\xi\in\mathbb{R}^{d}\,\mbox{ and for a.e. }y\in Y,$
for some $\mu>0$. We will be commenting more on the initial datum $u^{0}_{\varepsilon}$ later.
Our present work is inspired by
  • a result due to Grégoire Allaire and Rafael Orive [3] dealing with the parabolic problems with large drift and potential terms
  • an article of Grégoire Allaire and Anne-Lise Raphael [4] which dealt with parabolic problems with large drift and potential terms in periodic porous media
  • a work of Yves Capdeboscq [7] addressing the periodic homogenization of eigenvalue problems for elliptic operators involving large drift terms
We also recall a work due to Grégoire Allaire and co-authors [2] dealing with parabolic problems with large potential terms (and no drift terms). Our present work is likewise related to the work of Patrizia Donato and Andrey Piatnitski [10]. In this work, we demonstrate that the solution to the initial boundary value problem (1.1) can be factorized as follows:
$u_{\varepsilon}(t,x)=e^{\frac{-\lambda t}{\varepsilon^{2}}}v_{\varepsilon}(t,x)\psi\left(\frac{x}{\varepsilon}\right)$ (1.2)
where $(\lambda,\psi)$ are an eigenpair associated with a certain exponential $Y$-periodic eigenvalue problem. Factorization similar to (1.2) are already available in [4, 3, 10]. The use of exponential $Y$-periodic eigenvalue problems in homogenization can be traced back to [6, page 383]. Allaire and Orive in [3, Lemma 2.1] (see also [7, Lemma 3], where Capdeboscq treats operators with no lower order terms) show the existence of a unique $\theta\in\mathbb{R}^{d}$ such that the eigenvalue problems (with $A^{\ast}$ denoting the adjoint of $A$):
$\left\{\begin{aligned} -\operatorname{div}_{y}({A}(y)\nabla\psi(y))+b(y)\cdot\nabla_{y}\psi(y)+c(y)\psi(y)=\lambda\psi(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi(y)e^{-2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$. (1.3)
and
$\left\{\begin{aligned} -\operatorname{div}_{y}({A}^{\ast}(y)\nabla\psi^{\ast}(y))-\operatorname{div}_{y}(b(y)\psi^{\ast}(y))+c(y)\psi^{\ast}(y)=\lambda^{\ast}\psi^{\ast}(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi^{\ast}(y)e^{2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$. (1.4)
have the following properties:
  • Both the problems (1.3) and (1.4) have the same first eigenvalue which is real, positive and simple.
  • The first eigenfunctions of (1.3) and (1.4) can be chosen such that
    $\min_{Y}\{\psi,\psi^{\ast}\}\geq a>0\qquad\mbox{ and }\qquad\int_{Y}\psi(y)\psi^{\ast}(y)\,{\rm d}y=1.$ (1.5)
  • The function $\beta:Y\to\mathbb{R}^{d}$ defined by
    $\beta(y):=\psi(y)\psi^{\ast}(y)b(y)+\psi(y)A^{\ast}(y)\nabla_{y}\psi^{\ast}(y)-\psi^{\ast}(y)A(y)\nabla_{y}\psi(y)\quad\mbox{ for }y\in Y,$ (1.6)
    is divergence-free and is of zero average over $Y$, i.e.
    $\int_{Y}\beta(y)\,{\rm d}y=0.$
The factorization (1.2) implies that $u_{\varepsilon}(0,x)$ and $v_{\varepsilon}(0,x)$ are equal to each other up to a factor of the first eigenfunction $\psi\left(\frac{x}{\varepsilon}\right)$ mentioned above. Observe further that, for the factorization (1.2) to hold, the factor $v_{\varepsilon}$ should necessarily satisfy the following initial boundary value problem:
$\left\{\begin{aligned} \sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(\alpha\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)+\frac{1}{\varepsilon}\beta\left(\frac{x}{\varepsilon}\right)\cdot\nabla v_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{\varepsilon}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$. (1.7)
where the coefficient vector field $\beta$ is defined in (1.6) and the coefficients $\sigma$ and $\alpha$ are $Y$-periodic and are defined as follows:
$\sigma(y):=\psi(y)\psi^{\ast}(y)\quad\mbox{ and }\quad\alpha(y):=\sigma(y)A(y)\quad\mbox{ for a.e. }y\in Y.$ (1.8)
Remark $1.$
The coefficient $\sigma$ is $Y$-periodic because
$\sigma(y)=\psi(y)\psi^{\ast}(y)=\psi(y)e^{-2\pi\theta\cdot y}e^{2\pi\theta\cdot y}\psi^{\ast}(y)$
Remark 1.2.
We have made a choice of taking the initial datum for $v_{\varepsilon}$ in (1.7) to be $\varepsilon$-independent. The factorization (1.2), then, forces the initial datum for $u_{\varepsilon}$ in (1.1) to be of the form
$u_{\varepsilon}(0,x)=u^{0}_{\varepsilon}(x)=u_{0}(x)\psi\left(\frac{x}{\varepsilon}\right)$
This is sometimes referred to as the well prepared initial datum. Similar well preparedness assumptions on the data have been made in the literature [2, 3]. This choice results in strong convergence of the factor $v_{\varepsilon}$ to its homogenized limit. If one were to keep the data $u_{\varepsilon}(0,x)$ independent of the parameter $\varepsilon$, then the factorization forces the initial datum for $v_{\varepsilon}$ to be
$v_{\varepsilon}(0,x)=\frac{u_{0}(x)}{\psi\left(\frac{x}{\varepsilon}\right)}.$ (1.9)
It should be noted that the equation (1.7) can still be homogenized with (1.9) as the initial datum. The homogenized equation remains the same, but with the initial datum for the limit problem being
$\left(\int_{Y}\frac{1}{\psi(y)}\,{\rm d}y\right)u_{0}(x).$
All of this can be argued using the notion of two-scale convergence [1], for example. In this case, the convergence will be weak and hence cannot talk about rates.
Our main result is the following quantitative homogenization result for the factor $v_{\varepsilon}$ in (1.2) which solves the initial boundary value problem (1.7).
Theorem 1.3.
Let $v_{\varepsilon}$ be the solution to (1.7) and let $v_{0}$ be the solution to the corresponding homogenized problem (1.12). Then, there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
$\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}v_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|v_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|u_{0}\right\|_{L^{2}(\Omega)}\right\}.$ (1.10)
In the above result, there is an inherent assumption that the terms on the right hand side of the bound are finite. This essentially translates to the data being sufficient smooth to guarantee adequate regularity for the solution $v_{0}$ of the homogenized equation. The proof of the above theorem is inspired by the ideas in the work of Jun Geng and Zhongwei Shen [11] that dealt with $\mathrm{L}^{2}$ convergence rates in the context of periodic homogenization of parabolic problems.
Remark 1.4.
Our estimate (1.10) is not sharp as in [11]. In there, the authors have obtained an $\mathcal{O}(\varepsilon)$ rate. Upon deriving a $\mathcal{O}(\sqrt{\varepsilon})$ rate for $\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$, the authors employ a duality argument to improve the rate to $\mathcal{O}(\varepsilon)$. We emphasise here that the same strategy can be followed in our case as well. Thus, improving the rate in Theorem 1.3 to $\mathcal{O}(\sqrt{\varepsilon})$.
Before we comment further on the proof of Theorem 1.3, we return to the equation (1.7). As $\beta$ (defined by (1.6)) is periodic, divergence-free and is of zero average, there exists a periodic skew-symmetric matrix-valued function $y\mapsto B(y)$ such that (see [7, Lemma 1, p. 809])
$\beta(y)=-\operatorname{div}_{y}B(y)\qquad\mbox{ in }Y.$
Taking a $Y$-periodic matrix-valued function $y\mapsto M(y):=\alpha(y)+B(y)$, the evolution equation in (1.7) can be rewritten as
$\sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(M\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)=0\qquad\mbox{ in }(0,T)\times\Omega.$ (1.11)
Observe that $M$ is also elliptic:
$M(y)\xi\cdot\xi=\alpha(y)\xi\cdot\xi\geq b\left|\xi\right|^{2}\quad\forall\xi\in\mathbb{R}^{d}\,\mbox{ and for a.e. }y\in Y,$
where the ellipticity constant $b$ depends on the ellipticity constant $\mu$ of $A$ and the lower bound $a$ in (1.5). Using classical homogenization arguments (see [9, Theorem 11.4, p. 211]), we can show that
$\lim_{\varepsilon\to 0}\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}=0,$
where $v_{0}$ satisfies the following homogenized initial boundary value problem:
$\left\{\begin{aligned} \frac{\partial v_{0}}{\partial t}-\operatorname{div}\left(M_{h}\nabla v_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{0}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$. (1.12)
where $M_{h}$ is the homogenized matrix with constant entries given by
$M_{h}=\int_{Y}\left(M(y)+M(y)\nabla_{y}\chi(y)\right)\,{\rm d}y,$
with $\chi(y)=\left(\chi_{1}(y),\dots,\chi_{d}(y)\right)$, where $\chi_{j}\in\mathrm{H}^{1}_{\text{per}}(Y)$ solve the cell problems:
$\left\{\begin{aligned} \operatorname{div}_{y}\left(M(y)\nabla_{y}\left(\chi_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\chi_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$. (1.13)
for $j=1,\dots,d$. Theorem 1.3 quantifies the aforementioned strong convergence of $v_{\varepsilon}$ to $v_{0}$.
Remark 1.5.
Factorizations comparable to (1.2) have appeared before [10, 4, 3]. We would like to emphasise here that [10] treats the setting where both the $\varepsilon$-problem and the corresponding homogenized problem are posed in the full space, i.e. the spatial domain is the whole of $\mathbb{R}^{d}$. The authors of [4] too are working in a perforated medium built out of the whole of $\mathbb{R}^{d}$. This is inherently due to the effective non-zero drift present in the model. In [3], the authors pose the problem in a bounded spatial domain $\Omega$. However, they make a change of unknown which takes the domain to the whole space $\mathbb{R}^{d}$ and incidentally the homogenized equation obtained in [3, Theorem 3.1] is posed in $\mathbb{R}^{d}\times[0,T)$. To the best of our knowledge, this present work is the first time where parabolic problems with large drift are treated and both the $\varepsilon$-problem and the homogenized problems are posed in $\Omega_{T}$ with $\Omega$ being a bounded domain in $\mathbb{R}^{d}$.
The results of this paper also yields quantitative rates for the periodic homogenization of parabolic equations with nondivergence form elliptic part and with large lower order terms. We draw inspiration from the work of Wenjia Jing and Yiping Zhang [12] which derived quantitative rates for the periodic homogenization of nondivergence form elliptic equations with large drift terms. The key idea in [12] is to use the invariant measure associated with the underlying diffusion process. This is a well-known procedure (see [5, 6], also see [8, section 2.2]) to transform nondivergence form equations into divergence form. This present work deviates from the earlier works by not appealing to this invariant measure. We demonstrate that via exponential eigenvalue problems akin to (1.3)-(1.4), one can transform the parabolic equation with terms in nondivergence form into a parabolic equation with terms in divergence form. The initial-boundary value problem, for the unknown $\rho_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, that we examine is the following:
$\left\{\begin{aligned} \frac{\partial\rho_{\varepsilon}}{\partial t}-\mathcal{K}\left(\frac{x}{\varepsilon}\right):\nabla^{2}\rho_{\varepsilon}+\frac{1}{\varepsilon}q\left(\frac{x}{\varepsilon}\right)\cdot\nabla\rho_{\varepsilon}+\frac{1}{\varepsilon^{2}}r\left(\frac{x}{\varepsilon}\right)\rho_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ \rho_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ \rho_{\varepsilon}(0,x)&=\rho_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (1.14)
Here the scalar-valued function $r$, the vector-valued function $q$ and the symmetric matrix-valued function $\mathcal{K}$ are all assumed to be $Y$-periodic. The coefficients $q$ and $r$ are assumed to be bounded in $\mathrm{L}^{\infty}(Y)$. We further assume $\mathcal{K}$ to be elliptic and that it belongs to $\mathrm{W}^{1,s}(Y)$ for some $s>d$. Invoking Sobolev embedding, one gets $\mathcal{K}\in\mathrm{C}^{0,\tau}(Y)$ for some $\tau\in(0,1]$. Additionally, we shall assume that $\operatorname{div}_{y}\mathcal{K}\in\mathrm{L}^{\infty}(Y)$. Thanks to these assumptions, the evolution equation in (1.14) can be rewritten as
$\frac{\partial\rho_{\varepsilon}}{\partial t}-\operatorname{div}\left(\mathcal{K}\left(\frac{x}{\varepsilon}\right)\nabla\rho_{\varepsilon}\right)+\frac{1}{\varepsilon}\left(\operatorname{div}_{y}\mathcal{K}\left(\frac{x}{\varepsilon}\right)+q\left(\frac{x}{\varepsilon}\right)\right)\cdot\nabla\rho_{\varepsilon}+\frac{1}{\varepsilon^{2}}r\left(\frac{x}{\varepsilon}\right)\rho_{\varepsilon}=0$
This is similar to the evolution equation in (1.1). Hence, analogous to the factorization (1.2), we shall have that the solution to (1.14) can be factored as follows:
$\rho_{\varepsilon}(t,x)=e^{\frac{-\widehat{\lambda}t}{\varepsilon^{2}}}\varphi_{\varepsilon}(t,x)\widehat{\psi}\left(\frac{x}{\varepsilon}\right),$ (1.15)
where $\widehat{\lambda}$ and $\widehat{\psi}$ are an eigenpair associated with the exponential eigenvalue problems akin to (1.3)-(1.4), with the following choices of the coefficients:
$A(y)=A^{\ast}(y)=\mathcal{K}(y);\qquad b(y)=\operatorname{div}_{y}\mathcal{K}(y)+q(y);\qquad c(y)=r(y).$
Thus, Theorem 1.3 gives a quantitative rate of convergence for $\varphi_{\varepsilon}$ to the solution of the corresponding homogenized equation.
2 Quantitative rates in $\mathrm{L}^{2}(\Omega_{T})$
In this section, we prove a general quantitative homogenized result for the following parabolic problem:
$\left\{\begin{aligned} \zeta\left(\frac{x}{\varepsilon}\right)\frac{\partial f_{\varepsilon}}{\partial t}-\operatorname{div}\left(\Theta\left(\frac{x}{\varepsilon}\right)\nabla f_{\varepsilon}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{\varepsilon}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (2.1)
In the above problem, the coefficient $\zeta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ and is assumed to satisfy
$\int_{Y}\zeta(y)\,{\rm d}y=1.$ (2.2)
The matrix-valued coefficient $\Theta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ is assumed to be uniformly elliptic, i.e.
$\Theta(y)\xi\cdot\xi\geq\kappa\left|\xi\right|^{2}\quad\forall\xi\in\mathbb{R}^{d}\,\mbox{ and for a.e. }y\in Y,$
for some $\kappa>0$. Considering the solutions $\omega_{j}\in\mathrm{H}^{1}_{\text{per}}(Y)$ to the cell problems
$\left\{\begin{aligned} \operatorname{div}_{y}\left(\Theta(y)\nabla_{y}\left(\omega_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\zeta(y)\omega_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$. (2.3)
for $j=1,\dots,d$, we can define a homogenized matrix as follows:
$\Theta_{h}=\int_{Y}\left(\Theta(y)+\Theta(y)\nabla_{y}\omega(y)\right)\,{\rm d}y,$
with $\omega(y)=\left(\omega_{1}(y),\dots,\omega_{d}(y)\right)$. Next, we consider the homogenized problem:
$\left\{\begin{aligned} \frac{\partial f_{0}}{\partial t}-\operatorname{div}\left(\Theta_{h}\nabla f_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ f_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ f_{0}(0,x)&=f_{\text{in}}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (2.4)
Our main result (Theorem 1.3) is a special case of the following quantitative result corresponding to the periodic homogenization of (2.1).
Theorem 2.1.
Let $f_{\varepsilon}$ be the solution to (2.1) and let $f_{0}$ be the solution to the corresponding homogenized problem (2.4). Then, there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
$\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$ (2.5)
Our line of approach is inspired by the work of Geng and Shen [11]. We introduce the following notation: for $s>0$,
$\Omega_{s}:=\left\{x\in\Omega\,\mbox{ such that }\mbox{ dist}(x,\partial\Omega)<s\right\}.$
We fix a space cut-off function $\eta_{1}^{\varepsilon}\in\mathrm{C}_{c}^{\infty}(\Omega)$ such that $0\leq\eta_{1}^{\varepsilon}\leq 1$ and it further satisfies:
$\eta_{1}^{\varepsilon}=1\quad\mbox{ in }\quad\Omega\backslash\Omega_{4\sqrt{\varepsilon}},\qquad\eta_{1}^{\varepsilon}=0\quad\mbox{ in }\quad\Omega_{3\sqrt{\varepsilon}},\qquad\left|\nabla\eta_{1}^{\varepsilon}\right|\leq\frac{1}{\sqrt{\varepsilon}}\quad\mbox{ in }\quad\Omega.$
We also fix a time cut-off function $\eta_{2}^{\varepsilon}\in\mathrm{C}_{c}^{\infty}(0,T)$ such that $0\leq\eta_{2}^{\varepsilon}\leq 1$ and it further satisfies:
$\eta_{2}^{\varepsilon}=1\quad\mbox{ in }\,\left(8\varepsilon,T-8\varepsilon\right),\qquad\eta_{2}^{\varepsilon}=0\quad\mbox{ in }\,(0,4\varepsilon]\cup(T-4\varepsilon,T),\qquad\left|\partial_{t}\eta_{2}^{\varepsilon}\right|\leq\frac{1}{\varepsilon}\quad\mbox{ in }\,(0,T).$
Next, we consider a parabolic smoothing operator borrowed from [11, Section 3, p.2102]:
$S_{\varepsilon}(g)(t,x):=\frac{1}{\varepsilon^{d+2}}\int_{\mathbb{R}^{d+1}}g(t-s,x-y)\,\theta\left(\frac{s}{\varepsilon^{2}},\frac{y}{\varepsilon}\right)\,{\rm d}y\,{\rm d}s,$ (2.6)
where $\theta\in\mathrm{C}^{\infty}_{c}(\mathcal{K})$ is a fixed nonnegative function that satisfies
$\int_{\mathbb{R}^{d+1}}\theta(s,y)\,{\rm d}y\,{\rm d}s=1.$
Here the set $\mathcal{K}\subset\mathbb{R}^{d+1}$ is defined below:
$\mathcal{K}:=\left\{(t,x)\in\mathbb{R}\times\mathbb{R}^{d}\mbox{ such that }\left|t\right|+\left|x\right|^{2}\leq 1\right\}.$
We shall also define the following subset of $\Omega\times(0,T)$:
$\Omega_{T,\delta}:=\Big(\Omega_{\delta}\times(0,T)\Big)\cup\Big(\Omega\times(0,\delta^{2})\Big)\cup\Big(\Omega\times(T-\delta^{2},T)\Big)\qquad\mbox{ for }\delta>0.$
In the Appendix, we have gathered few useful properties of the smoothing operator $S_{\varepsilon}$ which play a crucial role in our proof. Using the cut-off functions and the smoothing operator mentioned above, we define $w_{\varepsilon}:\Omega_{T}\to\mathbb{R}$ as follows:
$w_{\varepsilon}:=f_{\varepsilon}-f_{0}-\varepsilon\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right),$ (2.7)
where $f_{\varepsilon}$ and $f_{0}$ are solutions to (2.1) and (2.4) respectively. Note that $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$ and that $w_{\varepsilon}(0,\cdot)\equiv 0$. We begin our analysis with the following result. In the remainder of this manuscript, we will sometimes be using the following shorthand notations:
$\zeta_{\varepsilon}(x):=\zeta\left(\frac{x}{\varepsilon}\right),\quad\Theta_{\varepsilon}(x):=\Theta\left(\frac{x}{\varepsilon}\right),\quad\omega_{\varepsilon}(x):=\omega\left(\frac{x}{\varepsilon}\right)$
Lemma 2.2.
Let $w_{\varepsilon}$ be defined by (2.7). Then for any $\psi\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$, we have
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (2.8)
$\quad-\varepsilon\int_{\Omega_{T}}\zeta_{\varepsilon}\omega_{\varepsilon}\cdot\partial_{t}\left[S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right]\psi+\int_{\Omega_{T}}\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left\{\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\}\cdot\nabla\psi$
$\quad-\varepsilon\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}\omega_{\varepsilon}\cdot\nabla\psi+\varepsilon\int_{\Omega_{T}}\phi_{kij}\left(\frac{x}{\varepsilon}\right)\frac{\partial}{\partial x_{k}}\bigg\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\bigg\}\frac{\partial\psi}{\partial x_{i}},$
where the summation convention over repeated indices $i,j,k\in\{1,\dots,d\}$ is used.
Here $\phi_{\varepsilon}(x)=\phi\left(\frac{x}{\varepsilon}\right)$ with $\phi$ being a $Y$-periodic third-order anti-symmetric tensor such that
$\operatorname{div}_{y}\phi(y)=\Theta(y)+\Theta(y)\nabla_{y}\omega(y)-\Theta_{h}\qquad\mbox{ in }Y.$ (2.9)
Proof.
A direct computation leads to
$\zeta_{\varepsilon}\partial_{t}w_{\varepsilon}=\zeta_{\varepsilon}\partial_{t}f_{\varepsilon}-\zeta_{\varepsilon}\partial_{t}f_{0}-\varepsilon\zeta_{\varepsilon}\omega_{\varepsilon}\cdot\partial_{t}\left\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right\}$
and
$\Theta_{\varepsilon}\nabla w_{\varepsilon}$ $=\Theta_{\varepsilon}\nabla f_{\varepsilon}-\Theta_{\varepsilon}\nabla f_{0}-\Theta_{\varepsilon}\nabla_{y}\omega\left(\frac{x}{\varepsilon}\right)S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)-\varepsilon\Theta_{\varepsilon}\nabla\left\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right\}\omega_{\varepsilon}$
$=\Theta_{\varepsilon}\nabla f_{\varepsilon}-\Theta_{h}\nabla f_{0}+\left(\Theta_{h}-\Theta_{\varepsilon}\right)\nabla f_{0}-\left(\Theta_{\varepsilon}+\Theta_{\varepsilon}\nabla_{y}\omega\left(\frac{x}{\varepsilon}\right)-\Theta_{h}\right)S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)$
$\quad-\left(\Theta_{h}-\Theta_{\varepsilon}\right)S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)-\varepsilon\Theta_{\varepsilon}\nabla\left\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right\}\omega_{\varepsilon}$
$=\Theta_{\varepsilon}\nabla f_{\varepsilon}-\Theta_{h}\nabla f_{0}+\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left(\nabla f_{0}-S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right)-\varepsilon\operatorname{div}\left(\phi_{\varepsilon}\right)S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)$
$\quad-\varepsilon\Theta_{\varepsilon}\nabla\left\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right\}\omega_{\varepsilon},$
where we have used the equation (2.9) in the last step. Hence, for every $\psi\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$, we have
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi=-\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi$ (2.10)
$\qquad-\varepsilon\int_{\Omega_{T}}\zeta_{\varepsilon}\omega_{\varepsilon}\cdot\partial_{t}\left[S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right]\psi+\int_{\Omega_{T}}\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left\{\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\}\cdot\nabla\psi$
$\qquad-\varepsilon\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}\omega_{\varepsilon}\cdot\nabla\psi-\varepsilon\int_{\Omega_{T}}\operatorname{div}\left(\phi_{\varepsilon}\right)\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}\cdot\nabla\psi,$
where we have used the fact that $f_{\varepsilon}$ and $f_{0}$ satisfy the evolution equations in (2.1) and (2.4), respectively. Owing to the anti-symmetric nature of the third-order tensor (i.e. $\phi_{kij}=-\phi_{ikj}$ for $1\leq i,j,k\leq d$), it follows that (see [13, Remark 3.1.3]) for a nice function $\psi$, we have
$\frac{\partial}{\partial x_{k}}\left(\phi_{kij}\right)\frac{\partial\psi}{\partial x_{i}}=\frac{\partial}{\partial x_{k}}\left(\phi_{kij}\frac{\partial\psi}{\partial x_{i}}\right).$
Thus, an integration by parts in the last term of the right hand side of (2.10) results in (2.8). ∎
Remark 2.3.
Observe that for any $\tau\in\mathrm{L}^{2}_{\textrm{per}}(Y)$ satisfying
$\int_{Y}\tau(y)\,{\rm d}y=0,$
there exists a $\varpi\in\mathrm{H}^{2}_{\text{per}}(Y)$ such that
$\Delta_{y}\varpi=\tau\quad\mbox{ in }Y.$
Taking $g=\nabla_{y}\varpi$, we deduce that for any $\kappa\in\mathrm{H}^{1}_{0}(\Omega)$,
$\int_{\Omega}\tau\left(\frac{x}{\varepsilon}\right)\kappa(x)\,{\rm d}x=\varepsilon\int_{\Omega}\operatorname{div}\left(g\left(\frac{x}{\varepsilon}\right)\right)\kappa(x)\,{\rm d}x=-\varepsilon\int_{\Omega}g\left(\frac{x}{\varepsilon}\right)\cdot\nabla\kappa(x)\,{\rm d}x.$
As Sobolev embedding implies that $g\in\mathrm{L}^{\frac{2d}{d-2}}(\Omega)$, we can get the following bound:
$\left|\int_{\Omega}\tau\left(\frac{x}{\varepsilon}\right)\kappa(x)\,{\rm d}x\right|\leq C\varepsilon\left\|\nabla\kappa\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}.$ (2.11)
In the next lemma, we obtain a bound for the left hand side of (2.8).
Lemma 2.4.
Let $w_{\varepsilon}$ be defined by (2.7). Then for any $\psi\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$, we have
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi$ (2.12)
$\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}\left\|\psi\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}$
Proof.
It follows from (2.8) that
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},\psi\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla\psi\leq\left|\int_{\Omega_{T}}(\zeta_{\varepsilon}-1)\partial_{t}f_{0}\psi\right|$
$\quad+\varepsilon\left|\int_{\Omega_{T}}\zeta_{\varepsilon}\omega_{\varepsilon}\cdot\partial_{t}\left[S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right]\psi\right|+\left|\int_{\Omega_{T}}\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left\{\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\}\cdot\nabla\psi\right|$
$\quad+\varepsilon\left|\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}\omega_{\varepsilon}\cdot\nabla\psi\right|+\varepsilon\left|\int_{\Omega_{T}}\phi_{kij}\left(\frac{x}{\varepsilon}\right)\frac{\partial}{\partial x_{k}}\bigg\{S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\bigg\}\frac{\partial\psi}{\partial x_{i}}\right|$
$\quad=:I_{1}+I_{2}+I_{3}+I_{4}+I_{5}$
Thanks to the assumption (2.2) on $\zeta$, invoking (2.11) from Remark 2.3 we arrive at the following bound:
$I_{1}\leq C\varepsilon\int_{0}^{T}$ $\left\|\nabla\left(\partial_{t}f_{0}\psi\right)\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}\leq C\varepsilon\left\{\int_{0}^{T}\left\|\psi\nabla\partial_{t}f_{0}\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}+\int_{0}^{T}\left\|\partial_{t}f_{0}\nabla\psi\right\|_{\mathrm{L}^{\frac{2d}{d+2}}(\Omega)}\right\}$
$\leq C\varepsilon\left\{\int_{0}^{T}\left\|\nabla\partial_{t}f_{0}\right\|_{\mathrm{L}^{d}(\Omega)}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega)}+\int_{0}^{T}\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{d}(\Omega)}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega)}\right\}$
$\leq C\varepsilon\left\{\int_{0}^{T}\left\|\nabla\partial_{t}f_{0}\right\|_{\mathrm{L}^{d}(\Omega)}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega)}+\int_{0}^{T}\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{d}(\Omega)}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega)}\right\}$
$\leq C\varepsilon\left\{\left(\int_{0}^{T}\left\|\nabla\partial_{t}f_{0}\right\|^{2}_{\mathrm{L}^{d}(\Omega)}\right)^{\frac{1}{2}}+\left(\int_{0}^{T}\left\|\partial_{t}f_{0}\right\|^{2}_{\mathrm{L}^{d}(\Omega)}\right)^{\frac{1}{2}}\right\}\left\|\psi\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}$
$\leq C\varepsilon\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;\mathrm{W}^{1,d}(\Omega))}\left\|\psi\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}.$
Note that we have employed the Hölder inequality in the third and the fifth step while the Poincaré inequality in $\mathrm{H}^{1}_{0}(\Omega)$ is used in the fourth step. In order to handle the $I_{2}$ term, we recall from (2.3) that the product $\zeta\omega$ averages to zero over the unit cell $Y$. Thus, Remark 2.3 yields the existence of a matrix-valued function $\mathcal{D}$ (need not be skew-symmetric) bounded in $\mathrm{L}^{\infty}(Y)$ (thanks to elliptic regularity) such that
$\varepsilon\int_{\Omega_{T}}\zeta_{\varepsilon}\omega_{\varepsilon}\cdot\partial_{t}\left\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\}\psi=\varepsilon^{2}\int_{\Omega_{T}}\frac{\partial}{\partial x_{i}}\left(\mathcal{D}_{ij}\left(\frac{x}{\varepsilon}\right)\right)\partial_{t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]\psi$
On the right hand side of the above expression, performing an integration by parts leads to
$\varepsilon^{2}\int_{\Omega_{T}}\mathcal{D}_{ij}\left(\frac{x}{\varepsilon}\right)\frac{\partial^{2}}{\partial x_{i}\partial t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]\psi-\varepsilon^{2}\int_{\Omega_{T}}\mathcal{D}_{ij}\left(\frac{x}{\varepsilon}\right)\partial_{t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]\frac{\partial\psi}{\partial x_{i}}$
Hence it follows that
$I_{2}$ $\leq\varepsilon^{2}\left|\int_{\Omega_{T}}\mathcal{D}_{ij}\left(\frac{x}{\varepsilon}\right)\frac{\partial^{2}}{\partial x_{i}\partial t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]\psi\right|+\varepsilon^{2}\left|\int_{\Omega_{T}}\mathcal{D}_{ij}\left(\frac{x}{\varepsilon}\right)\partial_{t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]\frac{\partial\psi}{\partial x_{i}}\right|$
$=:J_{1}+J_{2}$
Note that
$\partial_{t}\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]$ $=S_{\varepsilon}\left(\partial_{t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)$
$=S_{\varepsilon}\left(\partial_{t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+\frac{\partial}{\partial x_{j}}S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial t}\right)-S_{\varepsilon}\left(\frac{\partial}{\partial x_{j}}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial t}\right).$
Therefore, we have
$J_{2}$ $\leq C\varepsilon^{2}\Bigg\{\left\|S_{\varepsilon}\Big(\partial_{t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\left\|\nabla\left(S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial t}\right)\right)\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\hskip 170.71652pt+\left\|S_{\varepsilon}\left(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial t}\right)\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\sqrt{\varepsilon}}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Note that
$\frac{\partial^{2}}{\partial x_{i}\partial t}$ $\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial}{\partial x_{i}}\left(\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)$
$=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)-S_{\varepsilon}\left(\frac{\partial}{\partial x_{i}}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)$
Hence we have
$J_{1}$ $\leq C\varepsilon^{2}\Bigg\{\left\|S_{\varepsilon}\Big(\partial_{t}\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\left\|\nabla\left(S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\partial_{t}\nabla f_{0}\Big)\right)\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\hskip 184.9429pt+\left\|S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\partial_{t}\nabla f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon^{\frac{3}{2}}}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\partial_{t}\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\sqrt{\varepsilon}}\left\|\partial_{t}\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
The above obtained estimates on $J_{1}$ and $J_{2}$ together will result in the following bound for $I_{2}$:
$I_{2}\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{H}^{1}(\Omega_{T})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Note that the product $\eta_{1}^{\varepsilon}(x)\eta_{2}^{\varepsilon}(t)$ takes the value one on $\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}}$. Hence we have
$I_{3}$ $\leq\left|\int_{\Omega_{T,4\sqrt{\varepsilon}}}\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left\{\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\}\cdot\nabla\psi\right|+\left|\int_{\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}}}\left(\Theta_{h}-\Theta_{\varepsilon}\right)\left\{\nabla f_{0}-S_{\varepsilon}(\nabla f_{0})\right\}\cdot\nabla\psi\right|$
$\leq\left\|\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}$
$\hskip 170.71652pt+\left\|\nabla f_{0}-S_{\varepsilon}(\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}$
$=:J_{3}+J_{4}$
Invoking the Calderón’s extension theorem for $f_{0}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))\cap\mathrm{H}^{1}(0,T;\mathrm{L}^{2}(\Omega))$, we can find a $\widetilde{f_{0}}:\mathbb{R}^{d+1}\to\mathbb{R}$ such that
$\left\|\nabla^{2}\widetilde{f_{0}}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}+\left\|\partial_{t}\widetilde{f_{0}}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}\leq C\Big\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Big\}$
Using this, we can bound $J_{4}$ as follows:
$J_{4}$ $\leq\left\|\nabla\widetilde{f_{0}}-S_{\varepsilon}(\nabla\widetilde{f_{0}})\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon\Big\{\left\|\nabla^{2}\widetilde{f_{0}}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}+\left\|\partial_{t}\widetilde{f_{0}}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d+1})}\Big\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon\Big\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Big\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon\Big\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;W^{1,d}(\Omega))}\Big\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})},$
where we have employed the Hölder inequality in the last step. Using a property of $S_{\varepsilon}$, we arrive at the following bound:
$J_{3}$ $\leq C\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T,5\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$=C\left\{\left(\int_{0}^{T}\int_{\Omega_{5\sqrt{\varepsilon}}}\left|\nabla f_{0}\right|^{2}\right)^{\frac{1}{2}}+\left(\int_{0}^{25\varepsilon}\int_{\Omega}\left|\nabla f_{0}\right|^{2}\right)^{\frac{1}{2}}+\left(\int_{T-25\varepsilon}^{T}\int_{\Omega}\left|\nabla f_{0}\right|^{2}\right)^{\frac{1}{2}}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{\frac{1}{4}}\Bigg\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\sup_{25\varepsilon<t<T}\left(\frac{1}{5\sqrt{\varepsilon}}\int_{t-25\varepsilon}^{t}\int_{\Omega}\left|\nabla f_{0}\right|^{2}\right)^{\frac{1}{2}}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})},$
where we have used the fact [11, Remark 3.6, page 2106] that for any $g\in\mathrm{H}^{1}(\Omega)$,
$\left\|g\right\|_{\mathrm{L}^{2}(\Omega_{\delta})}\leq C\sqrt{\delta}\left\|g\right\|_{\mathrm{H}^{1}(\Omega)}$
and the following estimate [11, Remark 3.10, page 2108]:
$\sup_{25\varepsilon<t<T}\left(\frac{1}{5\sqrt{\varepsilon}}\int_{t-25\varepsilon}^{t}\int_{\Omega}\left|\nabla f_{0}\right|^{2}\right)^{\frac{1}{2}}\leq C\left\{\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}$
Putting together the above estimates on $J_{3}$ and $J_{4}$ results in the following bound on $I_{3}$:
$I_{3}\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}+\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{\text{in}}\right\|_{\mathrm{L}^{2}(\Omega)}\right\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$
To estimate $I_{4}$, note that
$\nabla\bigg\{S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\bigg\}=S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)+S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla^{2}f_{0}\Big).$ (2.13)
Hence $I_{4}$ can be rewritten as
$I_{4}$ $\leq\varepsilon\left|\int_{\Omega_{T}}\Theta_{\varepsilon}S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)\omega_{\varepsilon}\cdot\nabla\psi\right|+\varepsilon\left|\int_{\Omega_{T}}\Theta_{\varepsilon}S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla^{2}f_{0}\Big)\omega_{\varepsilon}\cdot\nabla\psi\right|$
$\leq C\varepsilon\Bigg\{\left\|S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T,5\sqrt{\varepsilon}})}+\left\|S_{\varepsilon}\Big(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla^{2}f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon\Bigg\{\frac{1}{\sqrt{\varepsilon}}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\left\|\nabla^{2}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Using the property (2.13), and arguing as we did in the case of $I_{4}$, we can arrive at the following bound for $I_{5}$:
$I_{5}\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;H^{2}(\Omega))}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Collecting all the bounds derived above, we arrive at (2.12). Hence the proof. ∎
We are now equipped to prove Theorem 2.1.
Proof of Theorem 2.1.
As $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$, we take $\psi=w_{\varepsilon}$ in the estimate (2.12) obtained in the previous lemma to arrive at the following inequality:
$\int_{0}^{T}$ $\left\langle\zeta_{\varepsilon}\partial_{t}w_{\varepsilon},w_{\varepsilon}\right\rangle_{\mathrm{H}^{-1}(\Omega)\times\mathrm{H}_{0}^{1}(\Omega)}+\int_{\Omega_{T}}\Theta_{\varepsilon}\nabla w_{\varepsilon}\cdot\nabla w_{\varepsilon}$
$\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}.$
Using the coercivity assumption on $\Theta_{\varepsilon}$, the left hand side of the above inequality can be bounded from below by
$\int_{0}^{T}\frac{{\rm d}}{{\rm d}t}\left\|\sqrt{\zeta_{\varepsilon}}w_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\kappa\left\|\nabla w_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}=\left\|\sqrt{\zeta_{\varepsilon}}w_{\varepsilon}(T,\cdot)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\kappa\left\|\nabla w_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}.$
This helps us deduce
$\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}f_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|f_{{\rm{in}}}\right\|_{L^{2}(\Omega)}\right\}.$
Thanks to Poincaré inequality, we have
$\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\leq C\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))}$
Using the definition (2.7) of $w_{\varepsilon}$, we have
$\left\|f_{\varepsilon}-f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\varepsilon\left\|\omega\left(\frac{x}{\varepsilon}\right)\cdot S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0}\right)\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq\left\|w_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\varepsilon\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Hence the estimate follows. ∎
Appendix
Lemma 2.5.
For any function $\phi\in\mathrm{L}^{2}(\Omega_{T})$ with $\operatorname{supp}{\phi}\subset\subset\Omega_{T}$, we have the following
  1. 1.
[…]
  2. 2.
[…]
where the constant $C$ depends upon dimension $d$.
Lemma 2.6.
For any function $f\in\mathrm{L}^{2}(\mathbb{R};\mathrm{H}^{2}(\mathbb{R}^{d}))$, we have the following
[…]
where the constant $C$ depends upon dimension $d$.
Lemma 2.7.
For $Y$-periodic function $g$ and $f\in\mathrm{L}^{p}(\Omega_{T})$ consider any subset $\omega\subset\subset\Omega$ and $0<s<\tau<T$ then we have the following
  1. 1.
[…]
  2. 2.
[…]
where $\omega(\varepsilon):=\{x\in\Omega:\operatorname{dist}(x,\omega)<\varepsilon\}$, $0<\varepsilon<\min\{s,\tau\}$ and the constant $C$ depends upon dimension $d$.